Grade 12th passMechanics

Calculate the ionic strength of a solution that is 0.20 mol kg-1 in KCl(aq) and 0.10 mol kg-1 inCuSO4(aq).

Ishan karki
8 Years agoGrade 12th pass

1 Answer

Arun
8 Years ago
Dear Ishan
 

The total ionic strength is a sum of the ionic strengths of each electrolyte

I = I(KCl) + I(CuSO4) = 1/2(1x1 + 1x1)(b/b0)(KCl) + 1/2(1x22+1x22)(b/b0)(CuSO4) = 0.1 + 4x0.2 = 0.9.

Note, that the total ionic strength can be also calculated by summing the product 1/2 zi2(bi/b0) for each individual ion